OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. x/x+3+2/5=0

OpenStudy (ash2326):

We have \[\frac{x}{x+3}+\frac{2}{5}=0\] \[\text{We need to find x}\] Are you here @Kaupkel?

OpenStudy (anonymous):

Yepp I'm here

OpenStudy (ash2326):

Let's subtract both sides by 2/5, we get \[\frac{x}{x+3}+\frac{2}{5}-\frac{2}{5}=0-\frac{2}{5}\] Now we get \[\frac{x}{x+3}=\frac{-2}{5}\]Let's multiply both sides by \(5( x+3)\) We have \[\frac{x}{x+3}\times 5(x+3)=\frac{-2}{5} \times 5(x+3)\] Now let's cancel the common terms, from both sides \[\frac{x}{\cancel{x+3}}\times 5\cancel{(x+3)}=\frac{-2}{\cancel5} \times \cancel 5(x+3)\] We get \[5x=-2(x+3)\] Now Let's simplify the right side \[5x=-2x-6\] Add 2x to both sides \[5x+2x=-2x-6+2x\] We get \[7x=-6\] Divide both sides by 7 we get \[x=\frac{-6}{7}\] Do you understand this?